# From Words to Numbers: Your Large Language Model Is Secretly A Capable Regressor When Given In-Context Examples

### v1

@@ 1 Introduction
Similar to previous work Garg et al. (2022), we use (synthetic) regression datasets. Synthetic regression datasets have the following advantages:
@@
Formally, let $\mathbb{D}_{n}$ be a dataset consisting of $n$ input-output examples: $\mathbb{D}_{n}=\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$, where $x_{i}\in\mathbb{R}^{d}$, with $1\leq d\leq 20$ typically. We have $y_{i}=f(x_{i})$, where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $y_{i}\in\mathbb{R}$. We do not put any restrictions on $f$ and study functions ranging from simple linear predictions (i.e., $f(x)=ax+b$) to more complex and highly non-linear functions (e.g., $f(x)=10x+sin(5\pi x)+cos(6\pi x)$). We study how well various models such as LLMs (e.g., GPT-4), traditional supervised models (e.g., Random Forest), and unsupervised baselines (e.g., random prediction) are capable of predicting $y_{n+1}$ when given access to $n$ input-output examples (i.e., $\mathbb{D}_{n}$) and $x_{n+1}$.
Our study shows that pre-trained large language models (LLMs) display a surprisingly good performance on various regression tasks. For example, in Figure 1, without any parameter update, Claude 3 approaches the performance of a Linear Regression model and largely outperforms other supervised methods such as Random Forest or Gradient Boosting on a randomly generated linear regression dataset, with one informative variable out of two.
@@ 2.1.2 Non-Linear Regression Datasets
While we create these datasets with different random seeds, resulting in different $\mathbb{D}_{n}$, making a particular $\mathbb{D}_{n}$ very unlikely to have been seen by the LLMs during training, it is still possible that they have seen different $\mathbb{D}_{n}$ originating from the same generator function $f$. In an attempt to mitigate this risk, we created 4 new non-linear datasets. We describe them in the Appendix B. For example, one of these functions is: $y(x)=10x+sin(5\pi x)+cos(6\pi x)$, where $x\sim Uniform(0,1)$
@@ 2.1.3 Regression With Non-Numerical Inputs
To further investigate whether the models are able to learn abstract tasks beyond those subsumed by numerical regression Razeghi et al. (2022), we design the following tasks. We (randomly) map symbols (i.e., characters) to numbers (e.g., $a\rightarrow 1$). We then randomly sample a subset of these symbols in order to keep the context size manageable and to not need a large number of examples. We map the symbols to a numerical value by sampling a weight vector $w\in\mathbb{R}^{d}$ and doing a dot product between it and the corresponding values of each symbol. We use lowercase ASCII letters as our symbols (i.e., a $\dots$ z). We randomly sample 5 symbols which will serve as our vocabulary. We include the pseudocode in Appendix B.11.
@@ 2.2 Models
We experiment with three types of models: (1) large language models such as GPT-4, (2) supervised models such as Random Forest, and (3) heuristic-based unsupervised models such as random sampling. All models have access to the same train data and are evaluated on the same test partition. They have access to an input dataset $\mathbb{D}_{n}$ and are asked to predict the $y_{n+1}$ corresponding to the $x_{n+1}$. The train partition is used for in-context exemplars for LLMs and supervised training for supervised methods. Due to budget constraints and the context size limitations of the LLMs, we round input values to two decimal places. [^3] We provide extra experiments without rounding in Appendix L to show that the strong results we observed are not an artifact of rounding. We repeat each experiment with different random seeds.
@@
We use a total of 10 traditional supervised models, available in most statistical learning packages. We use: Linear Regression (4 versions: no regularization, Ridge regularization, and Lasso Regularization, no regularization and with polynomial features), Multi-Layer Perceptron (6 versions, 3 versions with different widths (Hornik et al., 1989) and 3 versions with different depths), Random Forest, Bagging, Gradient Boosting, AdaBoost, SVM, KNN, Kernel Ridge, and Splines. Similar to the LLM case, we do not tune any hyperparameters and use the defaults available in sklearn. It is important to note that these supervised baselines are very strong: (1) many of them are the results of algorithms specifically designed for regression (e.g., Splines); (2) all perform parameter updates (unlike an LLM with ICL); and (3) the default hyperparameters, as set in widely-used statistical packages, have been refined over time to offer reliable and generally strong performance across a variety of scenarios.
@@
Additional details on the models are provided in Appendix C. We include results with additional models, such as the latest release of GPT-4 (gpt-4-2024-04-09) or Mixtral Mixture of Experts 8x22B in the Appendix F, where we present the average rank obtain by each model.
@@ 3 Large Language Models Can Do Linear Regression
We present extended results, encompassing a wider array of models and datasets, in Appendix D.
@@ 4.1 Friedman Benchmarks
We use the 3 synthetic regression benchmarks introduced by Friedman (1991). Below, we provide the definition of the Friedman #2 dataset, with complete definitions for all datasets available in Appendix B.
@@ 4.2 New Regression Datasets
For the next dataset (called Original #2), we draw inspiration from the datasets introduced by Friedman, but we modify the domain of $x$ and change the operands (e.g., ${}^{2}$rightarrow ${}^{4}$). We provide an example below:
@@
It is important to underscore that the primary goal of these novel datasets is not to construct inherently difficult challenges for the LLMs, but rather to minimize the probability of evaluating them on datasets they could have already seen during their training phase. We provide additional details on these datasets in Appendix B, along with additional datasets. For an in-depth analysis of potential data contamination concerns, including additional experiments conducted to address these issues, refer to Appendix N.
@@ 4.3 Discussion
Overall, our results show that LLMs with ICL are capable of performing non-linear regression. For example, Claude 3 Opus outperforms Gradient Boosting and KNN on all 5 datasets. We present extended results, encompassing a wider array of models and datasets, in Appendix E. We observed that LLMs struggle on the datasets generated with randomly initialized neural networks (e.g., Simple NN #1, Transformer #1), although they remain, generally, better than the unsupervised methods.
Due to space constraints, we included in Appendix K an analysis of the performance of LLMs on non-numerical regression datasets. We found that even in this regime, LLMs outperform the unsupervised baselines.
@@ 5 How Fast Do Large Language Models Adapt?
Borrowing from the Online Learning community, we empirically analyze how the cumulative regret (i.e., cumulative loss) grows with respect to the time step (number of examples in context) Orabona (2019). Ideally, a good model should, over time, approach the quality of decisions that the best fixed strategy, informed by hindsight, would have made. In other words, the cumulative regret should, ideally, grow sub-linearly over time. To empirically estimate how the regret grows, we fit 3 curves: (1) Linear Fit: $a*x+b$, (2) Sqrt Fit: $a*sqrt(x)+b$ and (3) Log fit: $a*log(x)+b$ [^7] The choice of linear, square root, and logarithmic fits is motivated by their common appearance in theoretical regret bounds within the online learning literature. We then use the $R^{2}$ coefficient to determine which curve fit is better. We show two qualitative plots in Figure 7. We summarize the results in Table 1, recording the curve fit with the highest $R^{2}$ coefficient for each model. Since simply picking the best curve fit according to the $R^{2}$ score might tell an incomplete story, we include additional plots in Appendix G, covering multiple models and all seven datasets. We draw the following observations. First, the performance of large language models improves with the number of examples, suggesting the mechanism at play is capable of effectively leveraging more data. Second, we remark that very capable LLMs, such as Claude 3 or GPT-4 can obtain sub-linear regret, meaning that the predictions made by the LLM approach the quality of decisions that the best algorithm would have made, leading to near-optimal performance in the long run.
@@
We remark that there are differences between our empirical analysis and online learning. Firstly, while online learning often focuses on establishing theoretical regret bounds, our approach is empirical, we only empirically show that the regret of certain LLMs grow sub-linearly by using curve fitting and $R^{2}$. To address potential concerns of overfitting and enhance the robustness of our empirical findings, we repeated the experiment 3 times and averaged the cumulative regret. Second, our results are only for finite (and relatively small) time steps, diverging from the online learning norm of analyzing behavior as $T$ approaches infinity. To provide further evidence that the results are not an artifact of small T, we performed the following experiment. We used GPT-4 and recorded its performance across multiple training dataset sizes, ranging from $20$ to $500$. We have observed that the performance of GPT-4 continues to improve as the number of in-context exemplars increases, suggesting that, our results are not an artifact of limited time steps. We include the associated plots in Appendix O.
@@ 6 Related Work
Another line of research, which is close to our work, is that of investigating what types of “functions” can be learned through in-context learning Garg et al. (2022); Zhang et al. (2023); Xing et al. (2024). Notably, all these works do not use pre-trained LLMs, but specifically train a transformer for the task. Garg et al. (2022) shows empirically that standard transformers can be trained from scratch to perform in-context learning of linear functions. That is, they ¡..¿. Guo et al. (2024) investigates more complex function classes. Wei et al. (2023) shows that larger language models are able to overcome their semantic priors when shown input-label mappings. Zhang et al. (2023) train transformers with a single linear self-attention layer to in-context learn linear regression tasks, showing that transformers are capable of obtaining a performance competitive with the best linear predictor. Bhattamishra et al. (2024) experiment with training various models to in-context learn boolean functions. Although not the main focus of their work, they also experiment with pre-trained models such as Llama 2 and GPT-4, showing that they obtain a performance similar to nearest-neighbor baselines for boolean functions.
@@ 9 Limitation
Third, the issue of data contamination poses a challenge, given the opaque nature of training datasets for many LLMs. We have taken several steps to mitigate this risk: (i) Our analysis spans multiple LLMs, reducing the likelihood of all models being contaminated in the same way; (ii) We evaluated models with multiple random seeds on newly introduced datasets (alongside known ones like Friedman #1). In this way, we diminish the chance that these models have been directly exposed to the exact datasets during training; (iii) We included results with Falcon 40B, whose training data is publicly available (please see Appendix N for more details). We acknowledge that these measures do not eliminate the potential for data contamination entirely.
@@ Appendix A Appendix Structure
In Appendix B we provide additional details of the datasets we used.
In Appendix C we provide additional details of the models we used.
In Appendix D, we provide additional experiments and results to complement Section 3: Large Language Models Can Do Linear Regression.
In Appendix E, we provide additional experiments and results to complement Section 4: Large Language Models Can Do Non-Linear Regression.
In Appendix F, we show the average ranks obtain by each model across different dataset types.
In Appendix G, we provide additional experiments and results to complement Section 5: How Fast Do Large Language Models Adapt?
In Appendix J we detail how LLMs provided justifications for their prediction.
In Appendix K we include another experiment: regression task with non-numerical inputs.
In Appendix L we analyze the effects of rounding.
In Appendix M we analyze whether the performance of LLMs is similar with KNNs or not.
In Appendix N we analyze whether the results we have seen could be the effect of data contamination.
In Appendix O we analyze whether the performance of LLMs plateaus after a given number of in-context examples or not.
In Appendix P we analyze the performance of LLMs whose backbone architecture is different from Transformers.
Appendix B Datasets
@@ B.9 Original # 5
B.10 Neural Network Induced
@@ B.11 Non-Numerical Regression
Appendix C Models
@@ C.1 LLM
GPT-4 (Achiam et al., 2023)
Not available
OpenAI
Chat GPT (Ouyang et al., 2022)
Not available
OpenAI
Claude 3 Opus (Anthropic, )
Not available
Anthropic
Claude 3 Sonnet (Anthropic, )
Not available
Anthropic
Gemini Pro (Team et al., 2023)
Not available
Google
Mistral Medium
Not Available
Mistral
Mixtral Mixture of Experts 8x7B (Jiang et al., 2024)
Available
Mistral
Mistral 7B (Jiang et al., 2023)
Available
Mistral
Llama2 70B (Touvron et al., 2023)
Available
Meta
Code Llama2 70B (Rozière et al., 2023)
Available
Meta
Yi 34B (AI et al., 2024)
Available
01.ai
Available
Databricks
GPT-4 (20240409) (Achiam et al., 2023)
Not available
OpenAI
GPT-3 Davinci (Brown et al., 2020)
Not available
OpenAI
GPT-3 Babbage (Brown et al., 2020)
Not available
OpenAI
Claude 3 Haiku (Anthropic, )
Not available
Anthropic
Claude ${}^{\dagger}$
Not available
Anthropic
Claude ${}^{\dagger}$
Not available
Anthropic
Claude ${}^{\ddagger}$
Not available
Anthropic
Cohere Command R ${}^{\circ}$
Available
Cohere
Mixtral Mixture of Experts 8x22B
Available
Mistral
Falcon 40B Almazrouei et al. (2023)
Available
TII
Falcon 40B Instruct Almazrouei et al. (2023)
Available
TII
RWKV v4 14B Peng et al. (2023)
Available
Various*
StripedHyena Nous 7B Poli et al. (2023b)
Available
TogetherAI
@@
GPT-4
gpt-4-0125-preview
Chat GPT
gpt-3.5-turbo-1106
GPT-4 (20240409)
gpt-4-turbo-2024-04-09
GPT-3 Davinci
davinci-002
GPT-3 Babbage
babbage-002
@@
Claude 3 Opus
anthropic/claude-3-opus
Claude 3 Sonnet
anthropic/claude-3-sonnet
Gemini Pro
google/gemini-pro
Mistral Medium
mistralai/mistral-medium
Claude 3 Haiku
anthropic/claude-3-haiku
Claude v2.1
anthropic/claude-2.1
Claude v2.0
anthropic/claude-2.0
Claude v1.2
anthropic/claude-1.2
Cohere Command R Plus
cohere/command-r-plus
StripedHyena Nous 7B
togethercomputer/stripedhyena-nous-7b
@@
Mixtral Mixture of Experts 8x7B
mistralai/Mixtral-8x7B-Instruct-v0.1
Mistral 7B
mistralai/Mistral-7B-Instruct-v0.1
Llama 70B
meta-llama/Llama-2-70b-chat-hf
Code Llama 70B
codellama/CodeLlama-70b-Instruct-hf
Yi 34B
01-ai/Yi-34B-Chat
Mistral 7B v2
mistralai/Mistral-7B-Instruct-v0.2
@@
DBRX
accounts/fireworks/models/dbrx-instruct
Mixtral 8x22B
accounts/fireworks/models/mixtral-8x22b

### v3

@@ 1 Introduction
Similar to previous work (Garg et al., 2022), we use (synthetic) regression datasets. Synthetic regression datasets have the following advantages:
@@
Formally, let $\mathbb{D}_{n}$ be a dataset consisting of $n$ input-output examples: $\mathbb{D}_{n}=\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$, where $x_{i}\in\mathbb{R}^{d}$, with $1\leq d\leq 20$ typically. We have $y_{i}=f(x_{i})$, where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $y_{i}\in\mathbb{R}$. We do not put any restrictions on $f$ and study functions ranging from simple linear predictions (i.e., $f(x)=ax+b$) to more complex and highly non-linear functions (e.g., $f(x)=x+10sin(\frac{5\pi x}{100})+10cos(\frac{6\pi x}{100})$). We study how well various models such as LLMs (e.g., GPT-4), traditional supervised models (e.g., Random Forest), and unsupervised baselines (e.g., random prediction) are capable of predicting $y_{n+1}$ when given access to $n$ input-output examples (i.e., $\mathbb{D}_{n}$) and $x_{n+1}$.
tOur study shows that pre-trained large language models (LLMs) display a surprisingly good performance on various regression tasks. For example, in Figure 1, without any parameter update, Claude 3 approaches the performance of a Linear Regression model and largely outperforms other supervised methods such as Random Forest or Gradient Boosting on a randomly generated linear regression dataset, with one informative variable out of two.
@@ 2.1.2 Non-Linear Regression Datasets
While we create these datasets with different random seeds, resulting in different $\mathbb{D}_{n}$, making a particular $\mathbb{D}_{n}$ very unlikely to have been seen by the LLMs during training, it is still possible that they have seen different $\mathbb{D}_{n}$ originating from the same generator function $f$. In an attempt to mitigate this risk, we created 5 new non-linear datasets. We describe them in the Appendix C. For example, one of these functions is: $y(x)=x+10sin(\frac{5\pi x}{100})+10cos(\frac{6\pi x}{100})$, where $x\sim Uniform(0,100)$ (plotted in Figure 5).
@@ 2.1.3 Regression With Non-Numerical Inputs
To further investigate whether the models are able to learn abstract tasks beyond those subsumed by numerical regression (Razeghi et al., 2022), we design the following tasks. We (randomly) map symbols (i.e., characters) to numbers (e.g., $a\rightarrow 1$). We then randomly sample a subset of these symbols in order to keep the context size manageable and to not need a large number of examples. We map the symbols to a numerical value by sampling a weight vector $w\in\mathbb{R}^{d}$ and doing a dot product between it and the corresponding values of each symbol. We use lowercase ASCII letters as our symbols (i.e., a $\dots$ z). We randomly sample 5 symbols which will serve as our vocabulary. We include the pseudocode in Appendix C.11.
@@ 2.2 Models
We experiment with three types of models: (1) large language models such as GPT-4, (2) supervised models such as Random Forest, and (3) heuristic-based unsupervised models such as random sampling. All models have access to the same train data and are evaluated on the same test partition. They have access to an input dataset $\mathbb{D}_{n}$ and are asked to predict the $y_{n+1}$ corresponding to the $x_{n+1}$. The train partition is used for in-context exemplars for LLMs and supervised training for supervised methods. Due to budget constraints and the context size limitations of the LLMs, we round input values to two decimal places. [^3] We provide extra experiments without rounding in Appendix N to show that the strong results we observed are not an artifact of rounding. We repeat each experiment with different random seeds.
@@
We use a total of 10 traditional supervised models, available in most statistical learning packages. We use: Linear Regression (4 versions: (i) no regularization, (ii) Ridge regularization, (iii) Lasso Regularization, and (iv) no regularization and with polynomial features), Multi-Layer Perceptron (6 versions, 3 versions with different widths (Cybenko, 1989; Hornik et al., 1989) and 3 versions with different depths (Lu et al., 2017)), Random Forest, Bagging, Gradient Boosting, AdaBoost, SVM, KNN, Kernel Ridge, and Splines. Similar to the LLM case, we do not tune any hyperparameters and use the defaults available in sklearn. It is important to note that these supervised baselines are very strong: (1) many of them are the results of algorithms specifically designed for regression (e.g., Splines); (2) all perform parameter updates (unlike an LLM with ICL); and (3) the default hyperparameters, as set in widely-used statistical packages, have been refined over time to offer reliable and generally strong performance across a variety of scenarios.
@@
Additional details on the models are provided in Appendix D. We include results with additional models, such as the latest release of GPT-4 at the time of submission (gpt-4-2024-04-09) or Mixtral Mixture of Experts 8x22B in Appendix G, where we present the average rank obtained by each model.
@@ 3 Large Language Models Can Do Linear Regression
We present extended results, encompassing a wider array of models and datasets, in Appendix E.
@@ 4.1 Friedman Benchmarks
We use the 3 synthetic regression benchmarks introduced by Friedman (1991). Below, we provide the definition of the Friedman #2 dataset, with complete definitions for all datasets available in Appendix C.
@@ 4.2 New Regression Datasets
For the next dataset (called Original #2), we draw inspiration from the datasets introduced by Friedman, but we modify the domain of $x$ and change the operands (e.g., 2 $\rightarrow$ 4). We provide an example below:
@@
It is important to underscore that the primary goal of these novel datasets is not to construct inherently difficult challenges for the LLMs, but rather to minimize the probability of evaluating them on datasets they could have already seen during their training phase. We provide additional details on these datasets in Appendix C, along with additional datasets. For an in-depth analysis of potential data contamination concerns, including additional experiments conducted to address these issues, please refer to Appendix P.
@@ 4.3 Discussion
Overall, our results show that LLMs with ICL are capable of performing non-linear regression. For example, Claude 3 Opus outperforms Gradient Boosting and KNN on all 5 datasets. We present extended results, encompassing a wider array of models and datasets, in Appendix F. We observed that LLMs struggle on the datasets generated with randomly initialized neural networks (e.g., Simple NN #1, Transformer #1), although they remain, generally, better than the unsupervised methods. Due to space constraints, we included in Appendix M an analysis of the performance of LLMs on non-numerical regression datasets. We found that even in this regime, LLMs outperform the unsupervised baselines.
@@ 5 How Fast Do Large Language Models Adapt?
Borrowing from the Online Learning community, we empirically analyze how the cumulative regret (i.e., cumulative loss) grows with respect to the time step (number of examples in context) (Orabona, 2019). Ideally, a good model should, over time, approach the quality of decisions that the best fixed strategy, informed by hindsight, would have made. In other words, the cumulative regret should, ideally, grow sub-linearly over time. To empirically estimate how the regret grows, we fit 3 curves: (1) Linear Fit: $a*x+b$, (2) Sqrt Fit: $a*sqrt(x)+b$ and (3) Log fit: $a*log(x)+b$ [^7] The choice of linear, square root, and logarithmic fits is motivated by their common appearance in theoretical regret bounds within the online learning literature. We then use the $R^{2}$ coefficient to determine which curve fit is better. We show two qualitative plots in Figure 7. We summarize the results in Table 1, recording the curve fit with the highest $R^{2}$ coefficient for each model. Since simply picking the best curve fit according to the $R^{2}$ score might tell an incomplete story, we include additional plots in Appendix H, covering multiple models and all seven datasets. We draw the following observations. First, the performance of large language models improves with the number of examples, suggesting the mechanism at play is capable of effectively leveraging more data. Second, we remark that very capable LLMs, such as Claude 3 or GPT-4 can obtain sub-linear regret, meaning that the predictions made by the LLM approach the quality of decisions that the best algorithm would have made, leading to near-optimal performance in the long run.
@@
We remark that there are differences between our empirical analysis and online learning. Firstly, while online learning often focuses on establishing theoretical regret bounds, our approach is empirical, we only empirically show that the regret of certain LLMs grow sub-linearly by using curve fitting and $R^{2}$. To address potential concerns of overfitting and enhance the robustness of our empirical findings, we repeated the experiment 3 times and averaged the cumulative regret. Second, our results are only for finite (and relatively small) time steps, diverging from the online learning norm of analyzing behavior as $T$ approaches infinity. To provide further evidence that the results are not an artifact of small T, we performed the following experiment. We used GPT-4 and recorded its performance across multiple training dataset sizes, ranging from $20$ to $500$. We have observed that the performance of GPT-4 continues to improve as the number of in-context exemplars increases, suggesting that, our results are not an artifact of limited time steps. We include the associated plots in Appendix Q.
@@ 6 Related Work
Many previous papers studied in-context learning (Min et al., 2022; Dai et al., 2023; Pan et al., 2023; Kossen et al., 2024), either from a more theoretical standpoint (von Oswald et al., 2022; Xie et al., 2022; Wies et al., 2023; Ahn et al., 2023; Vladymyrov et al., 2024) or from a more empirical one (Garg et al., 2022; Akyürek et al., 2023; Zhang et al., 2023). For a more general discussion of in-context learning, we refer the reader to Dong et al. (2023). Our work fits into the empirical camp, showing through experiments that large language models pre-trained for next-token prediction on web-scale datasets are capable of doing regression when shown input-output examples in their context, sometimes with performance rivalling that of supervised methods such as Gradient Boosting or Random Forests. Previous studies Garg et al. (2022); Bai et al. (2023); Guo et al. (2024); Bhattamishra et al. (2024) have explicitly demonstrated the capacity of transformers to be trained for in-context learning, enabling them to implement a wide range of algorithms when provided with relevant examples. For example, it has been shown that transformers can be trained to in-context learn linear functions, with performance similar to that of optimal least squares (Garg et al., 2022; Zhang et al., 2023). Bai et al. (2023) extended this, showing that transformers can implement a broad class of standard machine learning algorithms in context when pre-trained for it. Li et al. (2023) proved generalization bounds for ICL under specific conditions. Different from this line of work, we do not train any model specifically for in-context learning or regression. Instead, we simply use large, powerful models (e.g., Claude 3, GPT-3, Llama2) and empirically show that they are capable of performing linear and non-linear regression when given in-context examples, despite not being specifically trained for it.
@@ 9 Limitations
Third, the issue of data contamination poses a challenge, given the opaque nature of training datasets for many LLMs. We have taken several steps to mitigate this risk: (i) Our analysis spans multiple LLMs, reducing the likelihood of all models being contaminated in the same way; (ii) We evaluated models with multiple random seeds on newly introduced datasets (alongside known ones like Friedman #1). In this way, we diminish the chance that these models have been directly exposed to the exact datasets during training; (iii) We included results with Falcon 40B, whose training data is publicly available (please see Appendix P for more details). We acknowledge that these measures do not eliminate the potential for data contamination entirely.
@@ Appendix A Appendix Structure
In Appendix B we expand on the related work discussion.
In Appendix C we provide additional details of the datasets we used, including their underlying formulas.
In Appendix D we provide additional details of the models we used, together with their exact API code.
In Appendix E, we provide additional experiments and results to complement Section 3: Large Language Models Can Do Linear Regression.
In Appendix F, we provide additional experiments and results to complement Section 4: Large Language Models Can Do Non-Linear Regression.
In Appendix G, we show the average ranks obtain by each model across different dataset types.
In Appendix H, we provide additional experiments and results to complement Section 5: How Fast Do Large Language Models Adapt?
In Appendix I, we provide additional experiments and results on real-world datasets.
In Appendix J, we provide results with more models from the Claude family.
In Appendix K we detail the total costs of running all the large language models.
In Appendix L we detail how LLMs provided justifications for their prediction.
In Appendix M we include another experiment: regression task with non-numerical inputs.
In Appendix N we analyze the effects of rounding.
In Appendix O we analyze whether the performance of LLMs is similar with KNNs or not.
In Appendix P we analyze whether the results we have seen could be the effect of data contamination.
In Appendix Q we analyze whether the performance of LLMs plateaus after a given number of in-context examples or not.
In Appendix R we analyze the performance of LLMs whose backbone architecture is different from Transformers.
Appendix B Related Work (Expanded)
Another line of research, which is close to our work, is that of investigating what types of “functions” can be learned through in-context learning (Garg et al., 2022; Zhang et al., 2023; Xing et al., 2024). Notably, all these works do not use pre-trained LLMs, but specifically train a transformer for the task. Garg et al. (2022) shows empirically that standard transformers can be trained from scratch to perform in-context learning of linear functions. Guo et al. (2024) investigates more complex function classes. Wei et al. (2023) shows that larger language models are able to overcome their semantic priors when shown input-label mappings. Zhang et al. (2023) train transformers with a single linear self-attention layer to in-context learn linear regression tasks, showing that transformers are capable of obtaining a performance competitive with the best linear predictor. Bhattamishra et al. (2024) experiment with training various models to in-context learn boolean functions. Although not the main focus of their work, they also experiment with pre-trained models such as Llama 2 and GPT-4, showing that they obtain a performance similar to nearest-neighbor baselines for boolean functions. A post on the AI Alignment Forum (Pesut, ) also explored the types of models that GPT-3 can fit in-context. While their focus was primarily on classification, they did examine regression scenarios as well.
While in this work we focus on in-context learning, there are other previous works that designed specific neural network modules (or cells) for arithmetic computations (Trask et al., 2018; Heim et al., 2020; Mistry et al., 2022).
Appendix C Datasets
@@ C.9 Original # 5
C.10 Neural Network Induced
@@ C.11 Non-Numerical Regression
C.12 Real-World Datasets
We also experimented with several real-world datasets: (i) Liver Disorders (UCI, 1990), (ii) Real Estate Valuation (Yeh, 2018), (iii) Diabetes (Efron et al., 2004), (iv) Servo (Ulrich, 1993), (v) Movies (Ahmed, 2017). The corresponding results are presented in Section I.
Appendix D Models
@@ D.1 LLM
GPT-4 (Achiam et al., 2023) Not available OpenAI
Chat GPT (Ouyang et al., 2022) Not available OpenAI
Claude 3 Opus (Anthropic, ) Not available Anthropic
Claude 3 Sonnet (Anthropic, ) Not available Anthropic
Gemini Pro (Team et al., 2023) Not available Google
Mistral Medium Not Available Mistral
Mixtral Mixture of Experts 8x7B (Jiang et al., 2024) Available Mistral
Mistral 7B (Jiang et al., 2023) Available Mistral
Llama2 70B (Touvron et al., 2023) Available Meta
Code Llama2 70B (Rozière et al., 2023) Available Meta
Yi 34B (AI et al., 2024) Available 01.ai
DBRX# Available Databricks
GPT-4 (20240409) (Achiam et al., 2023) Not available OpenAI
GPT-3 Davinci (Brown et al., 2020) Not available OpenAI
GPT-3 Babbage (Brown et al., 2020) Not available OpenAI
Claude 3 Haiku (Anthropic, ) Not available Anthropic
Claude v2.1 Not available Anthropic
Claude v2.0 Not available Anthropic
Claude v1.2 Not available Anthropic
Cohere Command R Plus Available Cohere
Mixtral Mixture of Experts 8x22B Available Mistral
Falcon 40B Almazrouei et al. (2023) Available TII
Falcon 40B Instruct Almazrouei et al. (2023) Available TII
RWKV v4 14B Peng et al. (2023) Available Various*
StripedHyena Nous 7B Poli et al. (2023b) Available TogetherAI
@@
GPT-4 gpt-4-0125-preview
Chat GPT gpt-3.5-turbo-1106
GPT-4 (20240409) gpt-4-turbo-2024-04-09
GPT-3 Davinci davinci-002
GPT-3 Babbage babbage-002
@@
Claude 3 Opus anthropic/claude-3-opus
Claude 3 Sonnet anthropic/claude-3-sonnet
Gemini Pro google/gemini-pro
Mistral Medium mistralai/mistral-medium
Claude 3 Haiku anthropic/claude-3-haiku
Claude v2.1 anthropic/claude-2.1
Claude v2.0 anthropic/claude-2.0
Claude v1.2 anthropic/claude-1.2
Cohere Command R Plus cohere/command-r-plus
StripedHyena Nous 7B togethercomputer/stripedhyena-nous-7b
@@
Mixtral Mixture of Experts 8x7B mistralai/Mixtral-8x7B-Instruct-v0.1
Mistral 7B mistralai/Mistral-7B-Instruct-v0.1
Llama 70B meta-llama/Llama-2-70b-chat-hf
Code Llama 70B codellama/CodeLlama-70b-Instruct-hf
Yi 34B 01-ai/Yi-34B-Chat
Mistral 7B v2 mistralai/Mistral-7B-Instruct-v0.2
@@
DBRX accounts/fireworks/models/dbrx-instruct
Mixtral 8x22B accounts/fireworks/models/mixtral-8x22b